INTRACAVITY SATURATED ABSORPTION SPECTROSCOPY OF THE 1.6-MICRON MOLECULAR OVERTONE BAND

\underline{H. SASADA}, AND K. SUZUMURA, Department of Physics, Faculty of Science and Technology, Keio University, Yokohama 223, Japan.

Saturation spectroscopy of the 2ν_{3} band of methane was carried out using an optical cavity absorption cell and a 1.6-μm-grating feedback external-cavity semiconductor laser. An experimental setup was similar with the pioneering work ^{a} in the 1.5-μm region. Since the linewidth of the observed saturation dips was just 0.5 MHz, a high potential exists for the use of this technique for high resolution spectroscopy and double-resonance spectroscopy.

aM. de Labachelerie, K. Nakagawa, and M. Ohtsu, Opt. Lett. 20, 410 (1994).